OpenStudy (anonymous):

An equation is shown below: 5(2x - 3) = 5 Part A: How many solutions does this equation have? Part B: What are the solutions to this equation? Show your work.

OpenStudy (kidrah69):

Ok so first lets solve it :) do you know where to start?

OpenStudy (kidrah69):

Hint: Start with distributing.

OpenStudy (xapproachesinfinity):

see 5 in the left and right can you do something about it?

OpenStudy (anonymous):

i don't know how to do these at all

OpenStudy (xapproachesinfinity):

why?

OpenStudy (xapproachesinfinity):

you don't know is not good enough

OpenStudy (kidrah69):

Multiply 5 by whats in the paranthese's

OpenStudy (xapproachesinfinity):

before that just multiply by 1/5 both sides of the equation to eliminate 5

OpenStudy (kidrah69):

Well you can do that at the end too :/

OpenStudy (xapproachesinfinity):

\[5(2x-3)=5 \Longrightarrow \frac{1}{5}5(2x-3)=\frac{1}{5}5\] so 2x-3=1

OpenStudy (xapproachesinfinity):

yeah you can to do that a well no problem just good to eliminate things as soon as possible

OpenStudy (anonymous):

you did not read the problem correctly that's not what i need

OpenStudy (playgames1122):

part A 10x-15=5 10x = 5+15 x= 20/10 x=2